Grade 11Physical Chemistry

Does the formula which is given in your notes for determination of total number of hybrid orbitals apply to molecules like CH4, NH3, etc. ?

Pranjal Singh
12 Years agoGrade 11

1 Answer

Kamlesh Kumar
12 Years ago
applying this formula you can find out hybridisation in any molecule....

S =sum of valence electrons, if S is less than or equal to 8
S
then no. of hybrid orbitals=----
2 S
and if S is greater than 8 then no of hybrid orbitals =-----
8
8
e.g in CH4 S = 4+1 X 4=8 , so no of hybrid orbitals=----= 4 i.e Sp3
2

e.g in I3-- S= 3 X 7 + 1 = 22
22
NO. OF HYBRID ORBITALS =----- = 2 + 3=5 that is Sp3 d hybridisation
8


after dividing by 8 or 2 if there is some remainder then divide it by 2 and add like 2+3



thanks&regards
kamlesh kumar
askiitians faculty